# Autonomous Exploration and Semantic Updating of Large-Scale Indoor Environments with Mobile Robots

### v2

@@ Abstract
We introduce a new robotic system that enables a mobile robot to autonomously explore an unknown environment, build a semantic map of the environment, and subsequently update the semantic map to reflect environment changes, such as location changes of objects. Our system leverages a LiDAR scanner for 2D occupancy grid mapping and an RGB-D camera for object perception. We introduce a semantic map representation that combines a 2D occupancy grid map for geometry with a topological map for object semantics. This map representation enables us to effectively update the semantics by deleting or adding nodes to the topological map. Our system has been tested on a Fetch robot, semantically mapping a $93m\times 90m$ and a $9m\times 13m$ indoor environment and updating their semantic maps once objects are moved in the environments.  [^1] Project page with code is available at: https://irvlutd.github.io/SemanticMapping
@@ I Introduction
In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes. To our knowledge, this is the first system that can perform all these tasks jointly and fully autonomously on a physical robot. Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner. Second, given the 2D map, the system plans a trajectory for the robot to revisit the whole environment. Third, the robot follows the planned revisiting trajectory and navigates in the environment. During navigation, we apply an open-vocabulary object detection and an object segmentation model in real-time to images from the RGB-D camera on the robot. The detected objects are added to the topological map as nodes in a graph structure. Consequently, a semantic map of the environment can be constructed once the robot finishes the trajectory. Finally, to handle changes in the environment, the robot can visit the environment again and update the semantic map by removing or adding nodes to the map. Fig.  1 illustrates an overview of our system.
We validated our system in a large-scale $(\sim 8,500m^{2})$ and a medium-scale $(\sim 117m^{2})$ indoor environments, using a Fetch robot. The robot successfully explored, mapped, and updated its semantic map of the scene by detecting and adapting to changes such as moved furniture.
@@
  • End-to-End Autonomy: We integrate unseen environment exploration, mapping, revisiting, and semantic updating into one system that runs autonomously on a physical robot online.
  • Practical Semantic Mapping: Using open-vocabulary object detection and segmentation, we build a hybrid representation of semantic map that efficiently scales to larger spaces.
  • Continuous Environment Adaptation: Our system enables real-time semantic map updates to reflect dynamic changes in the environment.
  • Physical System validation: We deploy and evaluate our system on a physical robot across large and medium-scale indoor environments, demonstrating the above contributions.
@@ II-B Robot Exploration with Semantics
II-C Semantic Mapping without Robot Exploration
In robot exploration, an exploration algorithm needs to be designed to control the movement of a robot, which also determines the movement of the sensors on the robot such as LiDAR scanners or RGB-D cameras. Thus, the robot will have the ability to actively explore the environment. There are a large number of methods in the literature that focus on semantic mapping from offline data [24, 13, 25, 26, 27, 28]. This data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera. Therefore, these systems are not fully autonomous and depend on human control to explore and map the environments.
Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method. For example, SemanticFusion [13] combines ElasticFusion [29] for dense reconstruction with a neural network for semantic segmentation. DA-RNN [25] combines KinectFusion [9] and a recurrent neural network for video semantic segmentation. [12] uses Mask R-CNN [30] for instance segmentation to construct and update a semantic map in real-time using a mobile manipulator system. However, it requires a fully explored environment map as a prerequisite to begin with. Hydra [31] builds and optimizes a 3D scene graph hierarchically from simulated and real-world datasets. Khronos [32] constructs dense spatio-temporal maps that reflect the 3D scene and also account for some dynamic environment changes while testing their method using a physical robot system. However, the robots are controlled manually during this process. Recently, open-vocabulary object detectors such as GroundingDINO [33] have been utilized for semantic mapping. For example, ConceptGraphs [14] and its hierarchical version [34] build 3D scene graphs with open-vocabulary object detectors based on a hand-held scanned dataset of the indoor environment.
Although these works significantly advance in semantic mapping, they all share a common limitation, i.e., they are not fully autonomous. As shown in Table I, prior methods are based on manual control for environmental exploration. Some methods require pre-explored maps, and some do not perform semantic updating of environments. In contrast, our method initially builds the map via robot exploration. Consequently, the robot can use this map to revisit the environment in a planned and greedy manner, while continuously refining and updating the semantic information in the environment. This offers a fully autonomous approach to both mapping and semantic updating.
Method Autonomous Exploration Mapping with Robot Semantic Mapping Semantic Updating
@@ II-C Semantic Mapping without Robot Exploration
TABLE I: Comparison of methods for autonomous exploration, semantic mapping and updating (reflecting environment changes).
@@ III System
Our system consists of two phases. In the initial phase, the mobile robot explores the environment and builds the occupancy map incrementally. In the subsequent phase, the robot leverages the built occupancy map to accurately localize itself within the environment and proceeds to construct the semantic map. The motivation to separate the process into two stages, rather than performing simultaneously, is because of computational efficiency and robustness. Attempting to perform both tasks in a single stage could lead to resource constraints and reduced performance, especially for large-scale environments with a large number of objects. Additionally, in large environments, as the number of detected objects increases, the number of map corrections (like loop closures) applied to the detected objects increase, which slows down the mapping process. Evidently, this has also been mentioned in [32]. Therefore, our two-phase approach is more efficient for semantic exploration and mapping of large-scale environments. An overview of the system is presented in Fig. 2.
@@ III-A Autonomous Exploration and Map Building
We have developed an environment exploration method based on the popular frontier exploration module [35], using a dynamic search window to efficiently navigate to frontiers in large real-world environments. When using [35] in large scale environments, as the map size increases, far-away frontiers become insignificant due to high cost, thus leaving a part of the map unexplored. Moreover, tuning the cost parameters in [35] is relatively hard and not intuitive.
@@
Dynamic Window Frontier Exploration. We modified the frontier exploration module [35] by dynamically adjusting the robot’s search area between local and global regions [^2] Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration. This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment. The search window is adjusted dynamically based on the density of the frontiers available around the robot. If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius. Setting this parameter to a large value expands the search region to the entire map. This allows the robot to move progressively ahead, reducing redundant coverage of the environment. The exploration continues until the number of frontiers falls below a global threshold or after a fixed time has elapsed, at which point the exploration terminates and saves the built map. Our improved exploration algorithm is integrated with the GMapping SLAM module [36, 37]. The approach is depicted in Fig. 2(a). During this process, the robot exploration trajectory is recorded which is later used to obtain a traversal plan. The saved occupancy map is then utilized to localize the robot during the semantic map construction and update phases.
III-B Autonomous Construction of Semantic Map
We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the whole environment to construct the semantic map.
@@ III-B2 Object Detection and Segmentation
During the traversal, object detection and segmentation is performed in real time to construct the semantic map. For real-time object detection, the system leverages GroundingDINO [33] to detect objects in the environment. It takes in the current RGB image from the robot’s camera, identifies objects, and generates labels and bounding boxes for each detection. These bounding boxes are used as prompts for MobileSAM [38], a faster version of the Segment Anything Model (SAM) [39], to generate the object segmentation masks.
After obtaining object segments of the current observation and its corresponding depth image, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in camera frame by applying camera intrinsic parameters to the depth. These points are then transformed to the map frame of reference through the transformation matrices representing camera pose in robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and robot pose in map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$:
$\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}.$ (1)
Finally, the mean position of the point cloud is computed using $\mathbf{P}_{\text{mean}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{P}_{\text{map}}(i)$ for $m$ points in the point cloud as the 3D position of the corresponding object in the map frame.
III-B3 Semantic Map Representation and Construction
Our semantic map is a hybrid representation. At the lower level, we utilize a 2D occupancy grid map $M$ that captures the spatial information of the free space and obstacles in the environment. At the higher level, we construct a topological map $G_{T}(V)$, where $V=\{v_{1},v_{2},\dots,v_{n}\}$ are the nodes representing uniquely detected objects (Fig. 2(d)). Each node $v_{i}$ has the following attributes.
@@ III-B3 Semantic Map Representation and Construction
This representation facilitates fast object association and easy update of the nodes in the event of changes in the environment.
Construction. To construct the semantic map, we first load the occupancy map $M$ built in Section III-A and localize the robot using it. Then an empty topological map $G_{T}(V)$, $V=\{\}$ is initialized. Additionally, for each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{\text{category}\}}$ is defined. The robot then starts tracking the traversal trajectory generated in Section III-B1. Simultaneously, object detection and segmentation module is executed to detect objects in the field of view of the robot and compute their positions in the map $M$ frame of reference as illustrated in Fig. 4(a).
As depicted in Fig. 4(b), for every detected object in the robot field of view, its minimum distance to the existing nodes of the corresponding category is calculated. If this distance is small than $\delta_{\{\text{category}\}}$, then that particular object is considered to be already represented on the map and hence not added to the topological map $G_{T}$ again. If the minimum distance is greater than the threshold, then the object is added as a new node to $G_{T}$ as illustrated in Fig. 4(b). The result of this partial construction step is shown in Fig. 4(c). This process is continued until the robot completes the tracking of the traversal trajectory, with new nodes being added to $G_{T}$ after object association. In this way, the semantic map is fully constructed, combining the geometric structure of the occupancy map with the layered semantics of the topological map.
Figure 4: Adding objects to semantic map.
Updating. Once some changes are made to the environment, we update the semantic map to reflect these changes. This process is similar to the construction phase, but with additional validation steps to handle existing object removal, displacement, and new object addition. First, the constructed semantic map is loaded, which acts as a reference for the current state of the environment. As the robot tracks the traversal trajectory again, the detected objects in its current field of view are compared against those objects that are expected to be in the same field of view in the semantic map. If an expected object is no longer detected, it is assumed to have been removed or moved and the corresponding node is deleted from the map as shown in Fig. 5. Similarly, newly detected objects that do not match any existing nodes within the distance threshold $\delta_{\{\text{category}\}}$ are added as new nodes, representing new or relocated objects. This process updates the semantic map in real-time.
Figure 5: Removing objects from semantic map.

### v3

# A Modular Robotic System for Autonomous Exploration and Semantic Updating in Large-Scale Indoor Environments

@@ Abstract
We present a modular robotic system for autonomous exploration and semantic updating of large-scale unknown environments. Our approach enables a mobile robot to build, revisit, and update a hybrid semantic map that integrates a 2D occupancy grid for geometry with a topological graph for object semantics. Unlike prior methods that rely on manual teleoperation or precollected datasets, our two-phase approach achieves end-to-end autonomy: first, a modified frontier-based exploration algorithm with dynamic search windows constructs a geometric map; second, using a greedy trajectory planner, environments are revisited, and object semantics are updated using open-vocabulary object detection and segmentation. This modular system, compatible with any metric SLAM framework, supports continuous operation by efficiently updating the semantic graph to reflect short-term and long-term changes such as object relocation, removal, or addition. We validate the approach on a Fetch robot in real-world indoor environments of approximately 8,5008,500m2 and 117117m2, demonstrating robust and scalable semantic mapping and continuous adaptation, marking a fully autonomous integration of exploration, mapping, and semantic updating on a physical robot.  [^1] Project page with code is available at: https://irvlutd.github.io/SemanticMapping
@@ I Introduction
In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes. Our system represents a significant step forward in integrating autonomous exploration, long-term semantic mapping, and updating on a physical robot, addressing limitations of prior methods that rely on manual control or posed datasets. Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner. Second, given the 2D map, the system plans a trajectory for the robot to revisit the whole environment. Third, the robot follows the planned revisiting trajectory and navigates in the environment. During navigation, we apply an open-vocabulary object detection and an object segmentation model in real-time to images from the RGB-D camera on the robot. The detected objects are added to the topological map as nodes in a graph structure. Consequently, a semantic map of the environment can be constructed once the robot finishes the trajectory. Finally, to handle changes in the environment, the robot can visit the environment again and update the semantic map by removing or adding nodes to the map. Fig.  1 illustrates an overview of our system.
We validated our system in a large-scale 8,5008,500m2 and a medium-scale 117117m2 indoor environments, using a Fetch robot. The robot successfully explored, and updated the semantic map by detecting changes such as moved furniture.
@@
  • End-to-End Autonomy: We integrate unseen environment exploration, mapping, revisiting, and semantic updating into a modular system that operates fully autonomously on a physical robot in real-time.
  • Lightweight Object Association: An efficient association strategy that enables fast and scalable semantic map updates.
  • Short and Long-Term Semantic Updating: Our semantic mapping approach captures changes during the current visit as well as changes observed across multiple revisits, enabling the map to reflect both immediate and accumulated environmental updates.
  • Real-World Validation at Scale: We deploy and evaluate our system on a physical robot in large- and medium-scale indoor environments, demonstrating the above contributions.
@@ II-B Robot Exploration with Semantics
Method Autonomous Exploration Metric Mapping with Robot Semantic Mapping Semantic Updating
@@
TABLE I: Comparison of methods for autonomous exploration, metric mapping with a physical robot, semantic mapping and updating (reflecting environment changes).
II-C Semantic Mapping without Robot Exploration
In robot exploration, an exploration algorithm needs to be designed to control the movement of a robot, which also determines the movement of the sensors on the robot such as LiDAR scanners or RGB-D cameras. Thus, the robot will have the ability to actively explore the environment. There are a large number of methods in the literature that focus on semantic mapping of offline data [27, 13, 24, 28, 29, 30]. This data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera. Therefore, these systems are not fully autonomous and depend on human control to explore and map environments.
Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method. For example, SemanticFusion [13] combines ElasticFusion [31] for dense reconstruction with a neural network for semantic segmentation. DA-RNN [24] combines KinectFusion [9] and a recurrent neural network for video semantic segmentation. [12] uses Mask R-CNN [32] for instance segmentation to construct and update a semantic map in real-time using a mobile manipulator system. However, it requires a fully explored environment map as a prerequisite to begin with. Hydra [25] builds and optimizes a 3D scene graph hierarchically from simulated and real-world datasets. Khronos [26] constructs dense spatio-temporal maps that reflect the 3D scene and also account for some dynamic environment changes while testing their method using a physical robot system. However, robots are manually controlled during this process. Recently, open-vocabulary object detectors such as GroundingDINO [33] have been utilized for semantic mapping. For example, ConceptGraphs [14] and its hierarchical version [34] build 3D scene graphs with open-vocabulary object detectors based on a hand-h eld scanned dataset of the indoor environment, having ground truth camera poses.
Although these works significantly advance in semantic mapping, they all share a common limitation, i.e., they are not fully autonomous. As shown in Table I, prior methods are based on manual control for environmental exploration. Some methods require pre-explored maps, and some do not perform semantic updating of environments. In contrast, our method initially builds the map through robot exploration.
@@ II-C Semantic Mapping without Robot Exploration
Consequently, the robot can use this map to revisit the environment in a planned and greedy manner, while continuously refining and updating the semantic information in the environment. This offers a fully autonomous approach to both mapping and semantic updating.
@@ III System
Our system consists of two phases. In the initial phase, a mobile robot explores the environment and builds the occupancy map incrementally. In the second phase, the robot leverages the built occupancy map to accurately localize within the environment and performs semantic mapping. The motivation to separate the process into two stages, rather than performing simultaneously, is because of computational efficiency and robustness. Performing both tasks in a single stage leads to resource constraints and reduced performance, especially in large-scale environments with many objects. Additionally, in large environments, as the number of detected objects increases, the number of map corrections (like loop closures) applied to the detected objects increase, which slows down the mapping process. This has also been mentioned in [26]. Therefore, our two-phase approach is more suitable for semantic mapping of large-scale environments. Furthermore, this separation provides a significant advantage for Lidar-based mapping that relies on scan matching, such as GMapping [35] and Hector SLAM [36]. Robot poses are not reliably maintained during the initial mapping for these methods as they rely on scan-matching. Accurate robot poses are obtained during the localization within a built map. Our two-stage approach allows these techniques to benefit from a reliable robot pose and thus more reliable object poses during the semantic mapping stage. An overview of our system is presented in Fig. 2.
@@ III-A Autonomous Exploration and Map Building
We have developed an environment exploration method based on the popular frontier exploration module [37], using a dynamic search window to efficiently navigate to frontiers in large real-world environments. When using the original approach in large-scale environments, as the map size increases, far-away frontiers become insignificant due to high cost, thus leaving a part of the map unexplored. Moreover, tuning the cost parameters in the original method is not intuitive for large environments.
Dynamic Window Frontier Exploration (DWFE). We modified the original frontier exploration method  [37] by dynamically adjusting the robot’s search area between local and global regions, as detailed in Algorithm 1222Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration. This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment. The search window is adjusted dynamically based on the density of the frontiers available around the robot. If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius, as seen in Fig. 3. Setting this parameter to a large value expands the search region to the entire map. This allows the robot to move progressively ahead, covering the local regions well, reducing redundant coverage of the environment. The exploration continues until the number of frontiers falls below a global threshold or after a fixed time has elapsed, at which point the exploration terminates and saves the built map. Our improved exploration algorithm is integrated with the GMapping SLAM module [35, 38]. The approach is depicted in Fig. 2(a). During this process, the robot exploration trajectory is recorded, which is later used to obtain a traversal plan. The saved occupancy map is then utilized to localize the robot during the semantic map construction and update phases.
Algorithm 1 Dynamic Window Frontier Exploration
1:Input: Occupancy map $M$, local, global radii $r_{l}$, $r_{g}$, local, global threshold $T_{l}$, $T_{g}$, max time $T_{max}$, weights $\alpha,\beta$
2:Initialize robot position $p$, time $t\leftarrow 0$
3:while $t<T_{max}$ do
4:  $F\leftarrow$ frontiers within $r_{l}$ around $p$
5:  if $|F|<T_{l}$ then
6:   $F\leftarrow$ frontiers within $r_{g}$ around $p$
7:   if $|F|<T_{g}$ then
8:     break $\triangleright$ Terminate if not enough frontiers
9:   end if
10:  end if
11:  $f^{*}\leftarrow\arg\min_{f\in F}\left(\alpha\cdot distance(p,f)+\beta\cdot\text{size}(f)\right)$
12:  yield $f^{*}$ $\triangleright$ Output the best frontier for this iteration
13:  Navigate to $f^{*}$
14:  $t\leftarrow t+1$
15:end while
Figure 3: Our dynamic search window during exploration.
@@
III-B Autonomous Semantic Mapping
We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the environment again swiftly, to construct the semantic map.
@@ III-B2 Object Detection and Segmentation
During traversal, object detection and segmentation are performed in real time to build the semantic map. Our system utilizes GroundingDINO [33] to detect objects from the robot’s RGB image observation, providing labels and bounding boxes, which are then used as prompts for MobileSAM [39], a faster version of SAM [40], to generate segmentation masks.
After obtaining object segments, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in the camera frame by applying the intrinsic parameters of the camera to the depth image observation. These points are then transformed into the map frame $M$, through the transformation matrices representing the camera pose in the robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and the robot pose in the map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$, as follows: $\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}$. Finally, the mean position of the point cloud is computed and represented as the 3D position of the corresponding object in the map frame.
III-B3 Semantics Representation and Mapping
Our semantic map is a hybrid representation. At the lower level, we utilize a 2D occupancy grid map $M$ that captures the spatial information of the free space and obstacles in the environment. At the higher level, we construct a topological map defined by a set of vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$. Each vertex $v_{i}$ represents a uniquely detected objects (Fig. 2(d)). Each node $v_{i}$ has the following attributes.
@@ III-B3 Semantics Representation and Mapping
This representation facilitates fast object association and easy update of nodes during environment changes.
Semantic Mapping and Updating. To build the semantic map, we first load the occupancy map $M$ built as detailed in Sec. III-A and localize the robot within it. If the semantic mapping process is performed for the first time, a topological map is initialized with an empty set of vertex $V^{\prime}=\emptyset$. Otherwise, the previously constructed semantic map is loaded ( $V^{\prime}\leftarrow V$). For each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{\text{category}\}}$ is defined to determine whether a detected object corresponds to an existing one. Categories without explicit thresholds are assigned a default value. In practice, setting $\delta$ roughly equal to the typical object size improves the differentiation between nearby objects.
Algorithm 3 Semantic Mapping and Updating
1:Input: Traversal trajectory $T_{s}$, Distance thresholds $\Delta=\{\delta_{c}\mid c\in\text{categories}\}$
2:Output: Updated semantic map $G_{T}(V^{\prime})$
3:if no prior map exists then
4:  Initialize $V^{\prime}\leftarrow\emptyset$, id $\leftarrow 0$
5:else
6:  Load $G_{T}(V)$
7:  $V^{\prime}\leftarrow V$,  $\text{id}\leftarrow\max_{v\in V^{\prime}}v.\text{id}+1$
8:end if
9:while Traversing $T_{s}$ do
10:  $\mathcal{O}_{\text{current}}\leftarrow$ objects detected in FoV
11:  $\mathcal{V}_{\text{expected}}\leftarrow$ expected nodes in FoV from $G_{T}(V^{\prime})$
12:  $\mathcal{O}_{\text{unmatched}}\leftarrow\mathcal{O}_{\text{current}}$ $\triangleright$ Initialize unmatched objects
13:  for $v\in\mathcal{V}_{\text{expected}}$ do
14:   if no $o\in\mathcal{O}_{\text{current}}$ such that
[…]
17:     $V^{\prime}\leftarrow V^{\prime}\setminus\{v\}$ $\triangleright$ Remove node
18:   else
19:     $\mathcal{O}_{\text{unmatched}}\leftarrow\mathcal{O}_{\text{unmatched}}\setminus\{o\}$
20:   end if
21:  end for
22:  for $o\in\mathcal{O}_{\text{unmatched}}$ do
23:   $V_{c}\leftarrow\{v\in V^{\prime}\mid v.\text{category}=o.\text{category}\}$
24:   $d_{\min}\leftarrow\infty$
25:   if $V_{c}\neq\emptyset$ then
26:     $d_{\min}\leftarrow\min_{v\in V_{c}}\text{dist}(P_{\text{mean}}(o),v.\text{position})$
27:   end if
28:   if $d_{\min}>\delta_{o.\text{category}}$ then
29:     $v_{new}=(\text{id},o.\text{category},$
30:          $P_{\text{mean}}(o.\text{segment}),$
31:          $o.\text{confidence})$
32:     $V^{\prime}\leftarrow V^{\prime}\cup\{v_{new}\}$ $\triangleright$ Add new node to $V^{\prime}$
33:     $\text{id}\leftarrow\text{id}+1$
34:   end if
35:  end for
36:end while
37:Return $G_{T}(V^{\prime})$
Once initialized, the robot then begins following the traversal trajectory $T_{s}$ (see Sect. III-B1). Simultaneously, the object detection and segmentation module processes the current image observation of the robot’s field of view (FoV) to detect objects $\mathcal{O}_{\text{current}}$ and compute their positions in the map $M$ frame of reference as seen in Fig. 5(a). These detections are initially considered unmatched $\mathcal{O}_{\text{unmatched}}$.
At each step, an approximate polygon representing the field of view is derived from the limits of the current depth point cloud. Using this FoV polygon, the semantic map is queried to retrieve the nodes $\mathcal{V}_{\text{expected}}$, which represent the objects expected to be within the FoV. The Semantic Mapping procedure (described in Alg. 3 iteratively updates the vertex set $V^{\prime}$.
Node Removal. For each expected node $v\in\mathcal{V}_{\text{expected}}$, the algorithm checks for a matching object $o\in\mathcal{O}_{\text{current}}$ of the same category ($o.\text{category}=v.\text{category}$) and within the threshold distance $\text{dist}(P_{\text{mean}}(o),v.\text{position})\leq\delta_{v.\text{category}}$. If no match is found, the node $v$ is removed from $V^{\prime}$, indicating that the corresponding object is no longer present in the expected location (e.g., moved or removed). The matched objects are removed from $\mathcal{O}_{\text{unmatched}}$.
Node Addition. For each unmatched object $o\in\mathcal{O}_{\text{unmatched}}$, the algorithm identifies nodes $V_{c}=\{v\in V^{\prime}\mid v.\text{category}=o.\text{category}\}$ belonging to the same category. If no such category exists $V_{c}=\emptyset$ or an object is not associated with existing ones $\min_{v\in V_{c}}\text{dist}(P_{\text{mean}}(o),v.\text{position})>\delta_{o.\text{category}}$, a new node $v_{new}=(\text{id},o.\text{category},P_{\text{mean}}(o.\text{segment}),o.\text{confidence})$ is added to $V^{\prime}$. The object association mechanism is illustrated in Fig. 5(b). This iterative process continues until the robot completes the traversal of $T_{s}$. By continuously updating the vertex set $V^{\prime}$, the algorithm ensures that the semantic map captures both short-term and long-term environmental changes. The resulting map seamlessly integrates the geometric structure of the occupancy map with the semantic information of the topological map , enabling robust and faster scene understanding.
Figure 5: Updating Semantic Map through Object Association.